OpenStudy (anonymous):

What are the x-intercepts of the graph of the function g(x)=x²-4x-12?

OpenStudy (cwrw238):

x^2 - 4x - 12 = 0 solve for x

OpenStudy (anonymous):

do you divide by 2?

OpenStudy (anonymous):

or divide by 4 or i mean

OpenStudy (cwrw238):

no this can be factored

OpenStudy (cwrw238):

x^2 - 4x - 12 = x^2 - 6x + 2x - 12 can you continue?

OpenStudy (anonymous):

x²-4x-12?

OpenStudy (cwrw238):

no you need to factor this by grouping factor x^2 - 6x then 2x - 12

OpenStudy (cwrw238):

i'll do the first part for you x²-4x-12 = x^2 - 6x + 2x - 12 = x(x- 6) +

OpenStudy (anonymous):

then divide them? idk

OpenStudy (cwrw238):

now fact = 2x - 12

OpenStudy (cwrw238):

* factor

OpenStudy (anonymous):

x(2x-12)?

OpenStudy (cwrw238):

no - take the 2 out

OpenStudy (anonymous):

x(x-12)

OpenStudy (cwrw238):

no the common factor of 2x and 12 is 2 so its 2(x - 6)

OpenStudy (anonymous):

so it's (-6,0) and (2,0)?

OpenStudy (cwrw238):

so going back we have x(x - 6) + 2(x - 6) note that the 2 parenthese cantain the same thing

OpenStudy (cwrw238):

we can then simplify it to (x-6)(x + 2) now we put this equal to zero to find the intercepts on x-axis (x - 6)(x + 2) = 0 so x = 6 or x = -2 giving the points of interception (6,0) and (-2,0)

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

yeah, thank you